d...ex
Problem #15
Find the constant term that should be added to make the following expression a perfect square trinomial.
x^2+7x
Problem #22
Solve by using the quadratic formula
5x^2-4x+1=0
Problem #23
Solve by completing the square.
4x^2+2x-3=0

Answer by stanbon(75887)

Find the constant term that should be added to make the following expression a perfect square trinomial.
x^2+7x+(7/2)^2
--------------
Problem #22
Solve by using the quadratic formula
5x^2-4x+1=0
x=[4+-sqrt(4^2-4*5*1)]/10
x=[4+-sqrt(-4)]/10
x=[4+-2i]/10
x=(2/5)+(1/5)i or x=(2/5)-(1/5)i
---------------------------
Problem #23
Solve by completing the square.
4x^2+2x-3=0
4(x^2+(1/2)x-(3/4)=0
Divide both sides by 4 to gt:
x^2+(1/2)x+? = (3/4)+?
Complete the square:
x^2+(1/2)x+(1/4)^2 = (3/4)+(1/4)^2
Factor to get:
(x+(1/4))^2 = (13/16)
Take the square root to get:
(x+(1/4)) =+- (1/4)sqrt(13)
x=(1/4)+(1/4)sqrt(13) or x= (1/4)-(1/4)sqrt13
================
Cheers,
Stan H.